# SWE-Dev: Evaluating and Training Autonomous Feature-Driven Software Development

### v2

@@ Abstract
Large Language Models (LLMs) have shown strong capability in diverse software engineering tasks, e.g. code completion, bug fixing, and document generation. However, feature-driven development (FDD), a highly prevalent real-world task that involves developing new functionalities for large, existing codebases, remains underexplored. We therefore introduce SWE-Dev, the first large-scale dataset (with 14,000 training and 500 test samples) designed to evaluate and train autonomous coding systems on real-world feature development tasks. To ensure verifiable and diverse training, SWE-Dev uniquely provides all instances with a runnable environment and its developer-authored executable unit tests. This collection not only provides high-quality data for Supervised Fine-Tuning (SFT), but also enables Reinforcement Learning (RL) by delivering accurate reward signals from executable unit tests. Our extensive evaluations on SWE-Dev, covering 17 chatbot LLMs, 10 reasoning models, and 10 Multi-Agent Systems (MAS), reveal that FDD is a profoundly challenging frontier for current AI (e.g., Claude-3.7-Sonnet achieves only 22.45% Pass@3 on the hard test split). Crucially, we demonstrate that SWE-Dev serves as an effective platform for model improvement: fine-tuning on training set enabled a 7B model comparable to GPT-4o on hard split, underscoring the value of its high-quality training data. Code is available here https://github.com/DorothyDUUU/SWE-Dev.
@@ 1 Introduction
Figure 1: Overview of SWE-Dev, a software development dataset providing feature development tasks with feature description and codebase as input and test cases for evaluation. It is uniquely grounded in real-world repositories and paired with executable test suites, enabling reliable, functionally verifiable supervision. SWE-Dev is evaluated on 37 autonomous coding systems and supports advanced training paradigms like SFT, RL, and multi-agent training.
@@
Large Language Models (LLMs) are rapidly transforming autonomous programming, with capabilities extending from generating isolated code snippets to complex interactions within entire repositories [1, 2]. As LLMs increasingly engage at this repository scale, rigorously evaluating their proficiency in handling complex coding systems becomes paramount for guiding their advancement. Current prominent benchmarks, while valuable, still struggle to judge how well LLMs perform in realistic, end-to-end development settings (Table 1). For example, SWE-Bench [3] measures only localized bug fixes described by GitHub issues, and RepoBench [4] evaluates the completion of a few unrelated functions within a repository. However, they overlook the core tasks of developing and integrating significant new functionalities, which truly define how real-world codebases evolve.
The task of developing and integrating new functionalities is formally defined as feature-driven development (FDD) [5, 6], which consists of 40% coding tasks of all development efforts [7, 8]. FDD involves the end-to-end creation of new features, from interpreting requirements in large, existing codebases to generating functionally correct and integrated code (see Figure 1). FDD is how most modern software, from large applications to essential libraries, primarily evolves and delivers continuous value [9, 10]. Consequently, mastering FDD is a critical way towards achieving more comprehensive and genuinely autonomous programming capabilities with coding systems.
@@
Recognizing the central role of FDD and the limitations of current evaluation benchmarks, we introduce a feature-driven SoftWarE Development dataset, SWE-Dev, which is the first large-scale dataset designed to evaluate and train autonomous AI systems on real-world FDD tasks. It comprises 14,000 training and 500 test instances derived from over 1,000 open-source projects, and is distinguished by three key characteristics: (1) Realistic scale and complexity: SWE-Dev requires substantial code modifications (avg. 190 LOC across 3 files), challenging models with the cross-file dependencies, large contexts, and significant implementation scope characteristic of real-world feature development. (2) Robust and grounded evaluation: Each SWE-Dev sample is grounded in a real open-source repository, guided by a well-defined project requirement description (PRD), and evaluated using executable test cases to ensure the functional correctness of the proposed implementation. This design ensures alignment between task objectives and evaluation, enabling robust assessment and model supervision. (3) Verifiable training set with executable test suites: Uniquely, all 14,000 training instances are paired with runnable environments and executable unit tests, providing crucial execution-based feedback that enables effective Supervised Fine-Tuning (SFT) validation, Reinforcement Learning (RL) with accurate rewards, and Multi-Agent System (MAS) training, refer to Table 1.
Our extensive experiments using SWE-Dev reveal several critical insights. Firstly, Repository-level feature development is challenging: our findings show even top-tier models like Claude-3.7-Sonnet[11] and GPT-4o[12] solve only 22.45% hard samples and 68.70% easy samples with Pass@3. Secondly, MASs generally outperform single-agent baselines in modest margins. Interestingly, simple general-purpose multi-agent methods (e.g., Self-Refine[13], Reflexion[14]) often outperform more complex code-specific agents, while requiring fewer model calls and lower cost. Lastly, task-specific training on this task gets substantial gains on all training methods. After training, a 7B fine-tuned model is comparable to GPT-4o on hard subset after task-specific training.
These findings point to several promising directions for future research. First, the difficulty of FDD for LLMs necessitates enhancing LLMs’ core reasoning and long-context capabilities for software development. Second, current MAS designs often suffer from unnecessary communication overhead and limited coordination efficiency. Future work should explore lightweight agent architectures and better context-sharing mechanisms for repository-level development. Lastly, our initial experiments with RL and role-based multi-agent training show that training can be beneficial, but headroom remains. Future work could investigate multi-agent training and long-context RL with SWE-Dev.
@@
$\bullet$ We introduce SWE-Dev, the first real-world dataset for autonomous feature-driven software development. The dataset includes both training and test splits, each with runnable environments and test cases, enabling a wide range of evaluation and training.
$\bullet$ Our evaluations on SWE-Dev offer novel insights into the proficiency and deficiencies of various coding systems (chatbot LLMs, reasoning LLMs, and MAS) on complex FDD tasks.
$\bullet$ We demonstrate SWE-Dev enabling and validating diverse training paradigms (SFT, RL, and MAS training), establishing its utility for advancing training-based adaptation.
Table 1: Comparison of SWE-Dev with existing repository-level benchmarks. Task (FC: Function Completion, PG: Project Generation, LC: Line Completion, IS: Issue Solving), usage of real-repository, availability of training sets, Number of Samples, and task statistics are compared here. Detailed statistics information is demonstrated e.g., line of code (LOC), task description PRD length.
w. Testcases # Samples
Task Real Repo Train Existence Train Test Total w. Testcases Avg. PRD Tokens Avg. LOC
ComplexCodeEval[15] FC 7k 0 134.2 38.21
CoderEval[16] FC 234 234 119.26 20.64
DevEval[17] FC 2k 2k 91.5 112
rSDE-Bench[18] PG 53 53 1553 157.88
M2rc-Eval[19] LC 59k 0 0 1
RepoBench[4] FC 23k 0 0 89
SWE-Bench[3] IS 19k 2k 195.1 32.8
SWE-Dev FDD 14.5k 14.5k 1845.4 190
@@ 2.1 Coding benchmarks
LLMs show significant potential in coding tasks, driving the need for robust benchmarks. Early benchmarks such as HumanEval[20], MBPP[21], APPS[22], and CodeContests[23] primarily focus on isolated, function-level tasks. These benchmarks test for correctness in constrained settings: short snippets, well-specified inputs, and short expected outputs. While useful for early-stage capability testing, such tasks fall short of reflecting the complex, multi-file dependency and long contexts nature of real-world software development tasks. To address this, repository-level benchmarks emerged, such as SWE-Bench[3] (issue fixing), RepoBench[4], and M2RC-Eval[19] (code completion/understanding). Despite this progress, they often face two main issues: (1) The scope of required code generation or modification remains limited (e.g., avg. 32.8 LOC in SWE-Bench, 38.21 LOC in ComplexCodeEval[17]), inadequately simulating large-scale feature development or refactoring. (2) Weak or inconsistent evaluation protocols: several benchmarks [15, 19, 4] rely heavily on proxy metrics such as code similarity or static heuristics, which often fail to reflect functional correctness. This compromises both the robustness of evaluation and the comparability of results across models[24, 25]. SWE-Dev directly tackles these limitations by providing large-scale repository-level feature development tasks with executable unit tests. Its tasks involve substantial modifications, addressing shortcomings in both code scope and trainable environments, thereby significantly increasing task complexity and realism.
@@ 2.2 Code LLMs Training
Training LLMs for coding tasks typically involves three stages: pre-training, supervised fine-tuning (SFT), and reinforcement learning (RL). Pre-trained models such as StarCoder[24] and Phi[26] leverage massive code corpora to learn syntax and general programming patterns. To improve instruction following and task completion, many works adopt SFT. Code Alpaca[27] employs self-instruct generation, WizardCoder[28] leverages Evol-Instruct[29] to synthesize more complex instructions. However, SFT fundamentally lacks exploration: it teaches models to imitate ground-truth outputs rather than to reason or build autonomously[30]. Beyond SFT, RL frameworks such as CodeRL[31] utilize test-case-based feedback to optimize model behavior. While promising, both SFT and RL approaches largely focused on function-level tasks, limiting their applicability to more complex development scenarios. To address this, SWE-Gym[32] explores extending training to repository-scale tasks using multi-agent systems. However, due to the lack of an executable training set in SWE-Bench, SWE-Gym resorts to constructing a separate dataset of 2,438 tasks, ultimately yielding only 500 trajectory samples for training. In contrast, our proposed SWE-Dev provides a large-scale repository-level training set with runnable environments and unit-test-based supervision. It supports SFT, RL, and multi-agent training with executable feedback, enabling realistic and scalable development of code generation systems.
@@ 3 SWE-Dev
SWE-Dev is the first dataset designed to train and evaluate autonomous coding systems on feature-driven software development tasks. Each instance requires the model to implement a new capability within an existing codebase, based on a natural language requirement and evaluated through real-world unit tests. This section describes the construction of the dataset (§ 3.1), its core features (§ 3.2), and key statistics (§ 3.3).
@@ 3.1 Dataset Construction
To support realistic feature-level tasks and test-based evaluation, we begin by collecting real-world repositories that reflect common development scenarios. Specifically, we select 8,000 popular PyPI packages based on star counts. However, not all repositories are suitable: many lack usable tests or need sophisticated installation. Therefore, we applied a strict filtering process focused on test suite executability. Repositories were retained only if they met two criteria: (1) they include an identifiable test suite (e.g., using pytest or unittest), and (2) their test files could be executed successfully within the package Docker environment, with all tests passing. This ensures the resulting tasks are grounded in verifiable, runnable functionality. After filtering, we obtain 1,086 validated repositories (as of December 12, 2024) and 9,339 executable test files.
Step 2: Call Tree Generation
@@
To locate the specific functions and methods involved in implementing a feature, we capture the runtime interactions between test cases and their corresponding source code through dynamic analysis. This process has two main parts: (1) Dynamic analysis: We execute each test file using pytest inside a containerized Docker environment and apply Python’s built-in trace module to record all triggered functions in source code. This results multiple linear call chains that record the sequence of invoked source functions. (2) Call tree ensemble: We aggregate the call chains into into a hierarchical call tree, where the nodes of call tree represent functions, and edges capture dependency relationships. The call tree is rooted from test functions and followed by triggered source functions. The depth and width of the tree reflect the complexity of the feature logic, including nested structures and cross-file dependencies. These trees provide a precise mapping from test behavior to implementation code, enabling us to localize relevant functions and systematically control task difficulty later.
Step 3: Task Generation
@@
Once we have localized the implementation logic using call trees, we convert it into a feature development task by (1) masking the relevant code and (2) generating a natural language requirement for this feature. These two components constitute a typical development scenario in which a feature is functionally defined but not yet implemented. To achieve this, we perform the following: (1) Target function masking: We use structural properties of the call tree (e.g., depth and node count) to select function nodes that represent the core logic under test. The corresponding implementation code is removed from the repository, leaving a functional gap to fill. (2) Project Requirement Document (PRD) generation: We construct the feature description in PRD by using GPT-4o to synthesize a high-level module description from the test file and augmenting the masked function’s docstring with implementation-level details. These two elements are combined into PRD, which serves as the task prompt. See example in Fig. 9 and prompt in Appendix H.
@@ 3.2 Dataset Features
Controlled Complexity via Call Tree: Leveraging call-tree analysis, SWE-Dev enables systematic adjustment of task difficulty by adjusting dependency depth for task generation. This uniquely supports rigorous assessment of model capabilities against varying complexities, see § 8 discussion.
@@
Executable Training Support: SWE-Dev includes runnable environments and test cases for every sample, enabling training paradigms such as SFT and RL with execution-based feedback.
3.3 Statistics

### v3

# SWE-Dev: Evaluating and Training Autonomous End-to-End Feature-Driven Software Development

@@ Abstract
Large Language Models (LLMs) have shown strong capability in diverse software engineering tasks. However, feature-driven development, a highly prevalent real-world task that involves developing new functionalities for large, existing codebases, remains underexplored. We therefore introduce SWE-Dev, the first large-scale dataset (with 14,000 training and 500 test samples) designed to evaluate and train autonomous coding systems on real-world end-to-end feature-driven software development tasks. To ensure verifiable and diverse training, SWE-Dev uniquely provides all instances with a runnable environment and its developer-authored executable unit tests. This collection not only provides high-quality data for Supervised Fine-Tuning (SFT), but also enables Reinforcement Learning (RL) by delivering accurate reward signals from executable unit tests. We evaluated SWE-Dev across 17 base LLMs, 10 reasoning-focused LLMs, 10 multi-agent systems, and 8 tool-augmented LLM agents. Results show substantial headroom: the best single-turn model reaches only 22.51% Pass@1 on the hard split, while OpenHands agents improve to 56.44% but still leave many tasks unsolved. Code is available here https://github.com/DorothyDUUU/SWE-Dev.
@@ 1 Introduction
Figure 1: Overview of SWE-Dev, a software development dataset providing feature development tasks with feature description and codebase as input and test cases for evaluation. It is uniquely grounded in real-world repositories and paired with executable test suites, enabling reliable, functionally verifiable supervision. SWE-Dev is evaluated on 45 autonomous coding systems and supports advanced training paradigms like SFT, RL, and multi-agent training.
@@
Large Language Models (LLMs) are rapidly transforming autonomous programming, with capabilities extending from generating isolated code snippets to complex interactions within entire repositories [github_copilot, cursor_ide]. As LLMs increasingly engage at this repository scale, rigorously evaluating their proficiency in handling complex coding systems becomes paramount for guiding their advancement. Current prominent benchmarks, while valuable, still struggle to judge how well LLMs perform in realistic, end-to-end development settings (Table 1). For example, SWE-Bench [swebench] measures only localized bug fixes described by GitHub issues, and RepoBench [repobench] evaluates the completion of a few unrelated functions within a repository.
In contrast, real-world software development is dominated by feature-driven development, where developers design, implement, and integrate new functionality into large, existing codebases. Prior empirical studies show that feature development accounts for a substantial fraction of daily engineering work [harness2025getting_started_fdd, thum2014featureide]. Traditional feature development [coad1999java, palmer2002practical] typically involves intermediate processes such as developing an overall model or building feature lists. For coding systems, the critical way towards achieving more comprehensive and genuinely autonomous programming capabilities is on the final implementation of features. This end-to-end process, from interpreting requirements within large, existing codebases to generating functionally correct and integrated code, is what we refer to as end-to-end feature-driven software development.
@@
Recognizing the central role of feature development and the limitations of current evaluation benchmarks, we introduce an feature-driven SoftWarE Development dataset, SWE-Dev. which is the first large-scale dataset designed to evaluate and train autonomous AI systems on real-world end-to-end feature-driven software development tasks. It comprises 14,000 training and 500 test instances derived from over 1,000 open-source projects, and is distinguished by three key characteristics: ❶ Realistic scale and complexity: SWE-Dev requires substantial code modifications (avg. 190 lines of code across 3 files), challenging models with cross-file dependencies, long contexts, and feature-level implementation scope. ❷ Robust and grounded evaluation: Each sample is grounded in a real open-source repository, specified by a project requirement description (PRD), and evaluated using executable test cases to verify satisfaction of the intended implementation, ensuring reliable assessment and supervision. ❸ Verifiable training with executable test suites: All 14,000 training instances include runnable environments and executable unit tests, providing execution-based feedback that supports supervised fine-tuning (SFT), reinforcement learning (RL) with accurate rewards, and multi-agent system (MAS) training.
We validate SWE-Dev from both evaluation and training perspectives. For evaluation, we test (i) base LLMs (instruction-tuned and reasoning-focused), (ii) tool-augmented LLM agents implemented via OpenHands, and (iii) multi-agent LLM systems (MAS). The results show that feature development remains challenging: under direct single-turn generation, the best base model Claude-3.7-Sonnet achieves only 49.47% and 22.51% Pass@1 on the easy and hard split, with most models far below this level, indicating substantial headroom. In contrast, OpenHands agents substantially improve performance, reaching 56.44% on the hard split, but still leaving nearly half of tasks unsolved. Within our MAS suite, gains over a single baseline are modest (11.09% $\rightarrow$ 20.03% on hard), and simple refinement-based strategies can be more efficient than workflow-heavy designs. For training, SWE-Dev provides runnable tests and executable environments for every instance, enabling SFT and RL with execution feedback; we observe clear gains from test-supervised training, with a fine-tuned 7B model reaching performance comparable to GPT-4o on the hard split.
@@
❶We introduce SWE-Dev, the first real-world dataset for autonomous end-to-end feature-driven software development. The dataset includes both training and test splits, each with runnable environments and test cases, enabling a wide range of evaluation and training.
❷We conduct a comprehensive evaluation on SWE-Dev across LLMs (chatbot and reasoning-focused), tool-augmented LLM agents (via OpenHands), and multi-agent LLM systems, yielding insights into their capabilities and failure modes on feature development.
❸We demonstrate SWE-Dev enabling and validating diverse training paradigms (SFT, RL, and MAS training), establishing its utility for advancing training adaptation.
Table 1: Comparison of SWE-Dev with existing repository-level benchmarks. We compare task type, the average number of golden lines of code, whether using the real-world executable tests are available, and the number of executable environments.
Task type
Lines of
Code
Real
Tests
#Exec
Env
ComplexCodeEval Func Comp 38.2 0
CoderEval Func Comp 20.6 43
M2rc-Eval Line Comp 1.0 0
CrossCodeEval Line Comp 1.0 0
SWE-Bench(Train) Issue Solve 32.8 0
R2E-Gym Issue Solve 5.7 13
SWE-Smith Bug Fix 9.6 28
SWE-Synth Bug Fix 5.7 7
SWE-Dev Feature Dev 190.0 1061
@@ 2.1 Coding benchmarks
Early benchmarks such as HumanEval [humaneval], MBPP [mbpp], APPS [apps], and CodeContests [codecontest] mainly target isolated, function-level problems with short contexts and well-specified I/O, which is useful for measuring basic correctness but far from real software development. To better reflect repository-scale settings, a range of repository-level benchmarks have been proposed (Table 1), spanning function completion (e.g., ComplexCodeEval [complexcodeeval], CoderEval [codereval]), line-level completion/understanding (e.g., M2RC-Eval [m2rc-eval], CrossCodeEval [crosscodeeval]), and issue solving or bug fixing (e.g., SWE-Bench [swebench], R2E-Gym [r2e-gym], SWE-Smith [swe-smith], SWE-Synth [swe-synth]). Despite this progress, two limitations remain common: (1) the scope of required code changes is still small (often <40 LOC, e.g., 32.8 LOC in SWE-Bench and 38.2 LOC in ComplexCodeEval), which inadequately captures feature-scale development; and (2) evaluation is often weak or inconsistent, with many benchmarks lacking real executable tests or runnable environments, forcing reliance on proxy metrics that can miss functional correctness [starcoder, huang2022execution]. SWE-Dev directly addresses these gaps by providing large-scale, repository-level feature development tasks with substantial code modifications and real-world executable tests in runnable environments for every instance, substantially improving realism and enabling training with execution feedback.
@@ 2.2 Code LLMs Training
Training LLMs for coding tasks typically involves three stages: pre-training, supervised fine-tuning (SFT), and reinforcement learning (RL). Pre-trained models such as StarCoder [starcoder] and Phi [abdin2024phi4technicalreport] leverage massive code corpora to learn syntax and general programming patterns. To improve instruction following and task completion, many works adopt SFT. Code Alpaca [codealpaca] employs self-instruct generation, WizardCoder [wizardcoder] leverages Evol-Instruct [wizardlm] to synthesize more complex instructions. However, SFT fundamentally lacks exploration: it teaches models to imitate ground-truth outputs rather than to reason or build autonomously [intructgpt]. Beyond SFT, RL frameworks such as CodeRL [coderl] utilize test-case-based feedback to optimize model behavior. Early RL [rlhf] for code mostly targeted short, function-level problems with immediate execution rewards, but recent work has scaled RL to more realistic SWE settings: SWE-RL [swe-rl] applies policy-gradient updates in an agentless scaffold on SWE-Bench, while DeepSWE [deepswe] and SkyRL-v0 [skyrl] (see also [golubev2025training]) train interactive, multi-turn SWE agents over full repositories and tools. In parallel, SWE-Gym [swe_gym] explores repository-scale training via multi-agent systems; however, due to the lack of an executable training split in SWE-Bench, it constructs a separate dataset of 2,438 tasks and yields only 500 trajectory samples for training. In contrast, SWE-Dev provides a large-scale repository-level training set with runnable environments and unit-test-based supervision, enabling scalable SFT, RL, and multi-agent training with direct execution feedback. In contrast, our proposed SWE-Dev provides a large-scale repository-level training set with runnable environments and test-based supervision. It supports SFT, RL, and MAS training with executable feedback, enabling realistic and scalable development of code generation systems.
@@ 3 SWE-Dev
SWE-Dev is the first dataset designed to train and evaluate autonomous coding systems on feature-driven software development tasks. Each instance requires the model to implement a new capability within an existing codebase, based on a natural language requirement and evaluated through real-world unit tests. This section describes the construction of the dataset (§ 3.1), its core features (§ 3.2).
@@ 3.1 Dataset Construction
Step 1: Test File Collection To support realistic feature-level tasks and test-based evaluation, we begin by collecting real-world repositories that reflect common development scenarios. Specifically, we select 8,000 popular PyPI packages based on star counts. However, not all repositories are suitable: many lack usable tests or need sophisticated installation. Therefore, we applied a strict filtering process focused on test suite executability. Repositories were retained only if they met two criteria: (1) they include an identifiable test suite (e.g., using pytest or unittest), and (2) their test files could be executed successfully within the package Docker environment, with all tests passing. This ensures the resulting tasks are grounded in verifiable, runnable functionality. After filtering, we obtain 1,072 validated repositories (as of December 12, 2024) and 9,314 executable test files.
@@
Step 2: Call Tree Generation To locate the specific methods involved in implementing a feature, we capture the runtime interactions between test cases and their corresponding source code through dynamic analysis. This process has two main parts: (1) Dynamic analysis: We execute each test file using pytest inside a Docker environment and apply Python’s built-in trace module to record all triggered functions in source code. This results in multiple linear call chains that record the sequence of invoked source functions. (2) Call tree ensemble: We aggregate the call chains into a hierarchical call tree, where the nodes of call tree represent functions, and edges capture dependency relationships. The call tree is rooted from test functions and followed by triggered source functions. The depth and width of the tree reflect the complexity of the feature logic, including nested structures and cross-file dependencies. These trees provide a precise mapping from test behavior to implementation code, enabling us to localize relevant functions and systematically control task difficulty later.
@@
Step 3: Task Generation Once we have localized the implementation logic using call trees, we transform it into a feature development task by (1) masking the existing implementations of feature components and (2) generating a natural language requirement for this feature. These components represent a typical development scenario where a new feature needs to be added to an existing system, but its implementation is yet to be realized. To achieve this, we perform the following: (1) Target function masking: We use structural properties of the call tree (e.g., depth and node count) to identify key function nodes that represent the core logic for the feature. The corresponding implementation code is removed from the repository, creating a space for the new feature to be integrated. (2) Project Requirement Document (PRD) generation: We construct the feature description in PRD by using GPT-4o to synthesize a high-level module description from the test file and augmenting the masked function’s docstring with implementation-level details. These two elements are combined into PRD, which serves as the task prompt. See example in Fig. 7 and prompts in Appendix J.
@@ 3.2 Dataset Features
Controlled Complexity via Call Tree: Leveraging call-tree analysis, SWE-Dev enables systematic adjustment of task difficulty by varying the masked dependency depth during task generation, supporting rigorous assessment under controlled complexity levels (see § 5). We define test split Easy tasks as those with call-tree depth 1–2, and Hard tasks as those with depth $\geq 3$, which typically require implementing 5+ interdependent functions.
@@
Repository-Level Split for Clean Evaluation: SWE-Dev enforces a repository-level split between training and test sets to reduce contamination from shared repository context. The training split contains 1,033 repositories, while the test split contains 39 repositories, with no repository overlap between splits. More statistics are shown in § C.
4 Experiment
In this section, we empirically evaluate the effectiveness of various coding systems and training paradigms on SWE-Dev. We first compare the performance of single-LLM(§ 4.1) and MAS(§ 4.1.2) on the feature-driven software development tasks. Then, the effectiveness of different training approaches, including SFT (§ 4.2.1), RL (§ 4.2.2), and multi-agent training (§ 4.2.3) is discussed.
Setup. We employed the Pass@$k$ as an evaluation metric in SWE-Dev [humaneval]. For non-tool-use-agent inference code context, since SWE-Dev requires both the PRD and codebase as inputs. The codebases consist of many tokens (an average of 2̃0K lines, see Table 5), exceeding typical LLM context window. Thus, in all the experiments below, we provide only the relevant code context in the call tree rather than the entire codebase, since the call tree track exactly the code used for completion the task.
4.1 Testing Results
This section presents the performance of 17 chatbot LLMs, 10 reasoning LLMs, 10 multi-agent systems, and 8 tool-augmented LLM agents (via OpenHands) on SWE-Dev, under the single-LLM, multi-agent, and agentic settings. Full details of the evaluated methods are provided in Appendix H.1.
4.1.1 Single LLM Inference
SWE-Dev presents substantial challenges for current LLMs, revealing a clear gap between existing coding capabilities and real-world software engineering demands.
The left two part of Figure LABEL:fig:single_llm_performance reports Pass@1 performance of chatbot and reasoning LLMs on SWE-Dev (see full results in Table 7). We observe that: (1) LLMs perform better on the easy split than the hard split. (2) Performance generally scales with model size, especially for LLMs within the same family, aligning with our understanding of LLM capabilities. (3) Even the best-performing chatbot LLM: Claude-3.7-Sonnet [claude3.7]) achieves 19.74% on the Hard split and 53.09% on the Easy split. The reasoning LLM: Grok-3-beta achieves 53.63% and 18.97% on Easy and Hard split, respectively. This still falls short of achieving strong performance, indicating that current models are not yet fully capable of handling tasks that approximate the complexity of real-world development scenarios.
4.1.2 Multi-Agent Inference
Table D.2 compares the performance, call times and total costs of various MAS against the single-agent baseline driven by GPT-4o-mini. Details of MAS are given in AppendixH.1.
MAS generally outperforms single-agent baselines on complex tasks. While the single-agent approach achieves only 11.09% Pass@1 on hard tasks, Self Refine [self_refine] and EvoMAC [evomac] improve performance to 20.03% and 13.60%, respectively. These results highlight the advantage of MAS in solving complex, reasoning-intensive problems.
4.1.3 LLM Agent Inference
The right part in Figure LABEL:fig:single_llm_performance reports the performance of tool-augmented LLM agents evaluated on SWE-Dev using the OpenHands [wang2024openhands] framework (see Table 10 in Appendix D.3 for detailed number). Compared to direct single-LLM inference, agentic execution with tools and test feedback leads to substantial performance improvements.
[…]
Reasoning-enabled agents are strong under execution feedback. With OpenHands, reasoning-enabled (Thinking) agents achieve high performance on the Hard split (e.g. Kimi-K2-Thinking is comparable to Kimi-K2-Instruct (50.86% vs. 51.05% on Hard split), suggesting that the benefits of explicit reasoning are context-dependent and become most meaningful when coupled with execution signals.
Despite strong gains, feature development remains unsolved. Even with agentic scaffolding, a large fraction of tasks on the Hard split remain unsolved, highlighting the inherent difficulty of end-to-end feature development and leaving substantial headroom for future advances in agent design, planning, and long-horizon reasoning.
4.2 Training Results
In this section, we evaluate SWE-Dev’s support for different training methods, including SFT, RL. Additionally, we present preliminary results from our exploration of multi-agent training, offering an initial assessment of MAS-based learning. For detailed training setups, refer to the Appendix H.2.
4.2.1 Single LLM SFT
We conducted experiments on Qwen2.5-Instruct models of various sizes (0.5B, 1.5B, 3B, and 7B) to assess the impact of SFT on performance in SWE-Dev. Experimental setting is in Appendix H.3.
4.2.2 Single LLM RL
Pass@1 Pass@3
Method Easy Hard Easy Hard
vanilla 25.74 6.68 33.35 7.73
SFT 27.09 9.77 34.49 13.63
PPO (online RL) 28.30 12.25 32.69 14.33
DPO (offline RL) 25.89 10.36 31.32 14.66
SWE-Dev provides precise test cases enabling accurate rewards for coding tasks, supporting both online and offline RL. In this section, we explore the impact of RL on the Qwen2.5-7B-instruct using SWE-Dev. Considering the computational cost of RL, we limit our experiments in this section to 2k training samples. For full training setup, refer to the Appendix H.4.
Zero-shot SFT
Easy Hard Easy $\Delta$(%) Hard $\Delta$(%)
0.5B 6.39 1.00 12.12 +90 4.37 +337
1.5B 8.05 1.23 18.20 +126 7.64 +521
3B 15.93 5.27 27.53 +73 12.46 +136
7B 25.74 6.68 36.90 +43 18.89 +183
Both online and offline RL improve performance, especially on hard tasks. Table 2 shows that both PPO [ppo] and DPO [dpo] significantly improve Pass@1 performance, especially on the Hard split. Furthermore, PPO outperforms SFT (2k, aligned with RL) on the same training data. These findings highlight the advantages of RL training.
RL boosts one-shot success but not multi-sample gains. While RL fine-tuning yields improvements in Pass@1, it has minimal impact on Pass@3. Specifically, PPO achieves a Pass@1 of 28.30% on easy tasks, a noticeable increase from the base model’s 25.74%, but the Pass@3 remains lower than the SFT-training, even the original model’s performance. These results suggest that RL can be beneficial in refining Pass@1, particularly for more complex tasks, by increasing the model’s efficiency in generating correct answers in fewer attempts. However, this efficiency comes at the cost of reduced exploration. This aligns with findings from prior work [yue2025doesreinforcementlearningreally]. Therefore, while RL improves performance, significant headroom remains, and more advanced methods or further training are needed to achieve improvements across tasks.
Table 4: Comparison of multi-agent role-wise training, base MAS and single LLM’s performance on Qwen2.5-7B-Instruct. $\Delta$ indicates the relative improvement over the base MAS system. Partial fine-tuning of either agent also leads to consistent gains, demonstrating the effectiveness of role-specific supervision enabled by SWE-Dev. FT and base refers to finetuning and non-finetuning.
Org Coder Easy $\Delta$(%) Hard $\Delta$(%)
Single - - 25.74 - 6.68 -
MAS base base 26.64 - 7.39 -
FT base 30.04 +12.76 12.36 +67.25
base FT 31.42 +17.94 11.49 +55.48
FT FT 31.65 +18.80 12.70 +71.85
(a) Complexity analysis
(b) Metric Comparison
(c) PRD Quality Analysis
4.2.3 Multi-Agent Training
MAS has shown promising results on SWE-Dev, and we further investigate the training process of MAS on this dataset. As depicted in Fig. 5, the ground truth test case supervision in SWE-Dev enables EvoMAC [evomac] to improve its performance across multiple rounds of reasoning. This iterative refinement process motivates us to explore EvoMAC as the MAS for training in SWE-Dev. We apply rejection sampling to enhance agent performance via role-wise training.
5 Dataset Analysis
We analyze SWE-Dev’s task complexity, evaluation setup, and PRD quality to demonstrate its uniqueness and reliability. Analysis of Task Difficulty and Call Tree Characteristics. We analyze how task difficulty in SWE-Dev correlates with call tree complexity. As introduced in § 3.1, a call tree reflects the dynamic function invocation structure for this task, where nodes represent functions and edges denote their call relationships. We use two metrics: depth, indicating the maximum call nesting, and node count, representing the total number of distinct functions involved in the task. Fig. 4 shows that GPT-4o’s performance declines as depth and node count increase, revealing a strong correlation between structural complexity and task difficulty. This suggests that deeper and broader call structures introduce more functional requirements and interdependencies, making tasks more challenging.
Evaluation Method Precision. SWE-Dev uses execution-based evaluation with test cases, enabling precise performance signals. We compare metrics: Exact Match (EM) [m2rc-eval], Exact Sequence (ES) [m2rc-eval], CodeBLEU [codebleu], and Pass@3, using Qwen2.5 models and GPT-4o. As Fig 4 shows, Pass@3 best reflects capability scaling, separating models by size and quality. In contrast, EM, ES, and CodeBLEU show minimal variance, failing to distinguish models. This demonstrates that SWE-Dev’s test-case-based evaluation provides a more robust and realistic signal of model performance, better reflecting the functional correctness required in real-world software development.
PRD Quality. SWE-Dev includes a PRD for each task to simulate realistic developer-facing requirements, which are primarily derived from the original docstrings found within the repository source code. While many functions in open-source code include docstrings, we found that these are often incomplete—lacking clear descriptions of behavior, parameters, or edge cases. To improve instruction clarity without fabricating content, we lightly refine existing docstrings using GPT-4o, grounded in the related file and surrounding context. To evaluate instruction quality, we conducted a human assessment on 100 sampled tasks. Two experienced engineers rated the original and refined PRDs along Actionability, Completeness, and Clarity (Appendix E.2 includes human instruction). As shown in Fig. 4, refined PRDs consistently scored higher across all dimensions. This supports SWE-Dev’s goal of providing realistic, well-scoped requirements for reliable model evaluation.
6 Conclusion
In this work, we introduced SWE-Dev, the first dataset for evaluating and training autonomous coding systems on end-to-end feature-driven development task. SWE-Dev consists of 14,000 training and 500 test instances, each uniquely equipped with runnable environments and developer-authored executable unit tests, which provides essential execution-based feedback for advanced training paradigms like SFT, RL, and multi-agent learning. Our experiments show feature-driven software development is profoundly challenging for current autonomous coding systems. We also validate that training on SWE-Dev can yield encouraging performance gains. These findings validate SWE-Dev as a critical platform for identifying current limitations and fostering breakthroughs in AI-driven software development. We hope the release of SWE-Dev spurs innovation in long-context reasoning, agent orchestration, and execution-aware training towards genuinely autonomous software engineering.
Appendix A LLM Usage Disclosure
In our work, we utilized GPT-4o exclusively to enhance readability and improve language fluency during the writing process. We take full responsibility for the content of this publication, including any content generated by LLM.
Appendix B Other Related Work
RL for Software Engineering Earlier RL and RLHF work on code has mostly targeted short, function-level problems (e.g., HumanEval-style benchmarks), where the model produces a single solution in one shot with an immediate test-based reward. More recently, several works have applied RL to more complex SWE settings [golubev2025training]. SWE-RL [swe-rl] uses policy-gradient RL in an agentless scaffold on SWE-Bench, but still frames the task as single-turn patch generation, which sidesteps the challenges of stateful, multi-step interaction and long-horizon credit assignment. In contrast, DeepSWE [deepswe] and SkyRL-v0 [skyrl] scale RL to interactive, multi-turn SWE agents operating over full repositories, tools, and long contexts. Our work is complementary: rather than proposing a new agent or RL algorithm, SWE-Dev provides a large, PRD-driven, feature-level benchmark with executable rewards and examines how standard SFT and PPO behave on multi-file feature implementation. This makes SWE-Dev a natural substrate for future RL-based SWE agents in the spirit of these recent works.
Appendix C Dataset
C.1 Dataset Statistics